OpenStudy (anonymous):

i=sqrt(-1) A=(i^3)(i^29)(i^208)

OpenStudy (anonymous):

\[i^3\cdot i^{29}\cdot i^{208}=i^{3+29+208}=i^{240}\]Do you know anything about powers of i, and how they cycle?

OpenStudy (anonymous):

goes like \[i^0=1\] \[i^1=i\] \[i^2=-1\] \[i^3=-1\] \[i^4=1\] so take the integer remainder when you divide the exponent by 4, and you will get one of the above 4 possibilities

OpenStudy (anonymous):

since 4 divides 240 evenly you get 1

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw